TheSmartOne (thesmartone):

For fun (: The drawing messed up and the link doesn't work. The question was: \(\Large \frac{a}{b+c} + \frac{b}{a+c} = \frac{c}{a+b}=1\) Find \(\Large \frac{a^2}{b+c} + \frac{b^2}{a+c} + \frac{c^2}{a+b}=??\)

TheSmartOne (thesmartone):



TheSmartOne (thesmartone):

Link in case the drawing goes crazy in the future: https://i.imgsafe.org/7ec1a7e7ca.png

TheSmartOne (thesmartone):

@mayankdevnani let's do this :)

OpenStudy (mayankdevnani):

yw :D

OpenStudy (mayankdevnani):

i got it !

TheSmartOne (thesmartone):

mayankdevnani: hint:- multiply a number/variable with given expression

OpenStudy (mayankdevnani):

Another Hint :- Carefully observe both expressions, i mean compare their numerator and denominator part

TheSmartOne (thesmartone):

mhmmm, I'm just going to do something to get started let's call b + c = x a + c = y a + b = z so we have a/x + b/y + c/z = 1 and we're trying to find the value of a^2/x + b^2/y + c^2/z mhmmm, still stuck ;p

TheSmartOne (thesmartone):

mhmm, a different way...multiply it by a \(\Large a\left(\frac{a}{b+c} + \frac{b}{a+c} +\frac{c}{a+b}\right) = a(1)\) \(\Large \frac{a^2}{b+c} + \frac{ab}{a+c} +\frac{ac}{a+b} = a\) ummm... multiply it by b \(\Large \frac{ab}{b+c} + \frac{b^2}{a+c} +\frac{bc}{a+b} = b\) and multiply it by c \(\Large \frac{ac}{b+c} + \frac{bc}{a+c} +\frac{c^2}{a+b} = c\) now let's see... ummm... mhmmm

TheSmartOne (thesmartone):

oh, I meant multiply the original equation by a, and then original expression by b, and then the original expression by c, btw now we have three expressions, mhmmm.... ummmm

OpenStudy (mayankdevnani):

well going

OpenStudy (mayankdevnani):

why dun you add them ?

TheSmartOne (thesmartone):

I tried doing it mentally but there are the 3 fractions we want the a^2 + b^2 + c^2 and then the others have a different denominators so I got confused :P

OpenStudy (mayankdevnani):

add them show your work

jimthompson5910 (jim_thompson5910):

I found the interesting fact that a^3+b^3+c^3 = -abc but I don't know how useful it is

OpenStudy (mayankdevnani):

seems unimportant

TheSmartOne (thesmartone):

That's interesting Jim! :) Adding them: \[\frac{a^2}{b+c} + \frac{b^2}{a+c} +\frac{c^2}{a+b}+ \frac{ab}{a+c} +\frac{ab}{b+c} +\frac{ac}{a+b} +\frac{bc}{a+b}+\\ \frac{bc}{a+c}+ \frac{ac}{b+c} = a+b+c\]

jimthompson5910 (jim_thompson5910):

oh nvm, that's a much easier way that I didn't see

OpenStudy (mayankdevnani):

good tso,now take common factors out

TheSmartOne (thesmartone):

from the numerator or the denominator? :b

OpenStudy (mayankdevnani):

denominator

OpenStudy (dumbcow):

i get 0 , which makes the first condition seem impossible.

OpenStudy (mayankdevnani):

first condition seems impossible how ?

OpenStudy (dumbcow):

if 2nd expression is 0, this is only satisfied if a=b=c = 0 . Then you get 0/0 terms in 1st condition.

TheSmartOne (thesmartone):

\[\frac{a^2}{b+c} + \frac{b^2}{a+c} +\frac{c^2}{a+b} +\frac{ac}{a+b} +\frac{bc}{a+b}+\\ \frac{bc}{a+c}+ \frac{ab}{a+c} + \frac{ac}{b+c}+\frac{ab}{b+c} = a+b+c\] \[\frac{a^2}{b+c} + \frac{b^2}{a+c} +\frac{c^2}{a+b} +\frac{ac+bc}{a+b} + \frac{bc+ab}{a+c} + \frac{ac+ab}{b+c} = a+b+c\] OH I SEE IT NOW \[\frac{a^2}{b+c} + \frac{b^2}{a+c} +\frac{c^2}{a+b} +\frac{c\cancel{(a+b)}}{\cancel{a+b}} + \frac{b\cancel{(a+c)}}{\cancel{a+c}} + \frac{a\cancel{(b+c)}}{\cancel{b+c}} = a+b+c\] \[\frac{a^2}{b+c} + \frac{b^2}{a+c} +\frac{c^2}{a+b} +c+ b + a = a+b+c\] \[\frac{a^2}{b+c} + \frac{b^2}{a+c} +\frac{c^2}{a+b} =0\] The answer is 0!

TheSmartOne (thesmartone):

Well, they are adding up to 0 not multiplying to 0, mhmmm

OpenStudy (dumbcow):

oh nevermind, was ignoring denominators. a,b,c can still be negative

OpenStudy (mayankdevnani):

:D

TheSmartOne (thesmartone):

Thanks for the interesting question, Mayank. :D I might use it to test a few people in real life maybe xD

OpenStudy (alexandervonhumboldt2):

TSO that is not fun. That is a question for 8th grade.

OpenStudy (alexandervonhumboldt2):

such easy

OpenStudy (astrophysics):

Nice!

jaynator495 (jaynator495):